OpenStudy (anonymous):

((a^2-b^2)/(a^2-2ab+b^2) / (2a^2+5ab+3b^2)/(2a+3b))

OpenStudy (anonymous):

\[((a^2-b^2)/(a^2-2ab+b^2))/((2a^2+5ab+3b^2)/(2a+3b))\]

OpenStudy (anonymous):

Please help I think I know the answer, just want to confirm how I got it, so basically the steps

OpenStudy (anonymous):

ok give me time to write it down on here lol

OpenStudy (anonymous):

lol k

OpenStudy (anonymous):

add like terms so a^2/a^2 is a b^2 cancels out cause the positive and the negative. so the anser is (a-2ab) for that parenthesis now work on the next 2a^2/2a would be 2a and 3x^2/3b would be 3b so the answer to that is (2a+5ab+3b) now put them together, (a-2ab)/(2a+5ab+3b) is a/2a -3ab+3b understand? i'll write it as it looks k give me asec

OpenStudy (anonymous):

\[\frac{ a }{ 2a}-3ab+3b\]

OpenStudy (anonymous):

\[\frac{ a }{ 2a } -3ab +3b\]

OpenStudy (anonymous):



OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

yeah i think so

OpenStudy (anonymous):

thank you so much i think i gt it!

OpenStudy (anonymous):

welcome glad i could help :D

OpenStudy (anonymous):

actually is there anyway you could show me the steps I am alittle confused on some of them

OpenStudy (anonymous):

i need someone to show me the steos

OpenStudy (anonymous):

*steps

OpenStudy (anonymous):

nvm got it!!!!!!!!!!!!!!!

OpenStudy (anonymous):

sorry i can show you, ddnt mean to take so long was cooking dinner if you need help im here

OpenStudy (anonymous):

ok is there anyway u could show the steps just to make sure :) thanks so much!

OpenStudy (anonymous):

sure just give me time lol

OpenStudy (anonymous):

sure lol

OpenStudy (anonymous):

ok the first part are in parenthesis start there, ((a^2-b^2)/(a^2-2ab+b^2)) Now add like terms a^2/a^2 is a -b/b^2 is -b Now rewrite it (a-2ab-b) Now do the second part (2a^2+5ab+3b^2)/(2a+3b)) add like terms again 2a^2/2a is a 3b^2/3b is 3b Now rewrite it (a+5ab+3b) Now put the two together and divide (a-2ab-b)/(a+5ab+3b)=a+3ab+2b understand?

OpenStudy (anonymous):

yeas, so my final answer would be a/2a-3ab+3b

OpenStudy (anonymous):

I kinda just crossed out what was both on the top and bottom of each fraction is that ok?

OpenStudy (anonymous):

i dont believe so

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

cheer up, you can try it love